OpenStudy (anonymous):

Using a directrix of y = -2 and a focus of (1, 6), what quadratic function is created?

OpenStudy (anonymous):

A. f(x) = one eighth (x - 1)2 - 2 B. f(x) = -one eighth (x + 1)2 - 2 C f(x) = -one sixteenth (x + 1)2 - 2 D f(x) = one sixteenth (x - 1)2 + 2

OpenStudy (anonymous):

Ill give a medal and I'll fan you D:

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

D

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

thank you so kmuch